# A Transfer Learning-Based Method for Water Body Segmentation in Remote Sensing Imagery: A Case Study of the Zhada Tulin Area

### v1

@@ Abstract
To address the prevalent challenges of domain shift and small sample sizes in remote sensing image water body segmentation, this study proposes and validates a two-stage transfer learning strategy based on the SegFormer model. The approach begins by training a foundational segmentation model on a diverse source domain, where it achieves an Intersection over Union (IoU) of 68.80% on its validation set, followed by fine-tuning on data from the distinct target domain. Focusing on the Zhada Tulin area in Tibet—a region characterized by highly complex topography and spectral features—the experimental results demonstrate that this strategy significantly boosts the IoU for the water body segmentation task from 25.50% (for direct transfer) to 64.84%. This not only effectively resolves the model performance degradation caused by domain discrepancy but also provides an effective technical paradigm for high-precision thematic information extraction in data-scarce and environmentally unique remote sensing scenarios.
@@ 1 Introduction
The accurate acquisition of spatial information on water bodies, as one of the most crucial natural elements on the Earth’s surface, is of great significance for water resource management, ecological environment protection, and disaster risk assessment. In recent years, with the continuous development of remote sensing technology, high-resolution satellite imagery has provided powerful support for large-scale water body monitoring. However, traditional water extraction methods, such as the Normalized Difference Water Index (NDWI) [22, 39], the Automated Water Extraction Index (AWEI) [7], threshold segmentation [24], and active contour models [14], often struggle to achieve satisfactory results in complex surface environments, especially in areas where the spectral characteristics of water bodies are similar to the background and the terrain is complex. These traditional methods often require assistance from other indices, such as the Normalized Difference Vegetation Index (NDVI) [29] and the Soil-Adjusted Vegetation Index (SAVI) [13], to differentiate water from other features.
The rise of deep learning technology has brought new opportunities for water body extraction from remote sensing imagery [10, 46, 21, 43]. Convolutional Neural Networks (CNNs) [16], with their powerful feature learning capabilities, have shown excellent performance in various remote sensing applications. Early Fully Convolutional Networks (FCNs) [19] laid the foundation for pixel-level semantic segmentation. Subsequently, foundational architectures like U-Net [28], SegNet [1], and the DeepLab series [3, 4] significantly improved segmentation accuracy by integrating techniques like atrous (dilated) convolutions [41] and powerful backbones like ResNet [11]. More recently, Transformer-based models like SegFormer [38], leveraging the global modeling capabilities of the attention mechanism [34] and advantages in multi-scale feature fusion, have shown particularly outstanding performance in semantic segmentation tasks. Hierarchical vision transformers like the Swin Transformer [18] have further advanced the field. These methods can automatically learn complex spatio-spectral features, significantly enhancing the accuracy and robustness of water body identification.
However, the success of deep learning models heavily relies on a large amount of high-quality annotated data. In the remote sensing domain, this requirement faces practical challenges: the acquisition of high-quality imagery is constrained by various factors, and precise pixel-level annotation is not only time-consuming but also requires extensive expertise. More critically, imaging conditions and feature characteristics vary significantly across different geographical regions, causing models trained in one area to experience a sharp performance decline when directly applied to another—a phenomenon widely known as "domain shift" in computer vision.
To mitigate these issues, transfer learning offers an effective pathway [25, 40]. By transferring knowledge learned from a data-rich source domain to a data-scarce target domain, it is possible to achieve better model performance with limited annotated samples. In the field of natural image processing, the "pre-train and fine-tune" paradigm based on ImageNet pre-training has become a standard practice [15]. However, remote sensing images possess unique spectral properties, spatial resolutions, and feature distribution patterns, making the simple application of transfer learning strategies from natural images often yield limited results [32]. How to design effective transfer learning methods tailored to the characteristics of remote sensing applications remains a question worthy of in-depth exploration. In recent years, researchers have explored various deep learning methods for remote sensing image analysis, including multi-attention networks [17], building extraction with sparse tokens [2], Local Attention Embedding Network (LANet) [44], and hybrid Transformer-CNN networks [42], all of which have achieved good results in different applications. Meanwhile, methods like non-local neural networks [37] and scene classification frameworks [31] have also provided new insights for remote sensing image understanding.
The Zhada Tulin in the Ali Prefecture of Tibet is a unique geomorphological landscape on the Qinghai-Tibet Plateau, and its associated gully water systems exhibit typical characteristics of a high-altitude arid region. Due to special geological structures and climatic conditions, water bodies in this area are characterized by small scales, irregular shapes, and significant seasonality. Furthermore, their spectral features are heavily influenced by sediment content, making them easily confused with the surrounding earth forest landscape. These factors pose a great challenge for water body extraction in the Zhada area, making it an ideal testbed for evaluating the performance of water segmentation algorithms. At the same time, remote sensing data for this region is relatively scarce, and labeled samples are limited, which aligns with the "small-sample" scenario often encountered in practical applications.
@@
Against this background, this study designs a two-stage transfer learning strategy for remote sensing water body segmentation. We first construct a large-scale source domain dataset containing diverse geographical environments to train a base model with good generalization capabilities. Then, we use a small amount of labeled data from the Zhada Tulin area to fine-tune the model, enabling it to adapt to the specific water body characteristics of the target region. Through this "general-to-specific" learning strategy, we expect to achieve high-precision water body segmentation under small-sample conditions. This research not only provides technical support for water information extraction in the Zhada region but also offers a valuable exploration for the application of transfer learning in remote sensing thematic information extraction.
@@ 2.1 Study Area and Datasets
The datasets for this study consist of a diverse source domain (Dataset A) and a highly challenging target domain (Dataset B).
@@
The source domain (Dataset A) was constructed to provide rich and generalizable learning samples for pre-training the foundational model. The imagery for this dataset is primarily sourced from China’s Gaofen-7 (GF-7) satellite, whose sub-meter high resolution (0.8-meter panchromatic and 3.2-meter multispectral) provides an excellent data foundation for fine-grained feature recognition. The spatial coverage of the source domain is extensive, focusing on diverse plateau and mountainous environments. It not only includes several typical regions in western China (such as arid lakes in Xinjiang, salt lakes in Qinghai, and plateau lakes in Tibet) but also encompasses mountain lakes and glacial meltwater rivers in the Rocky Mountains of the United States. The variety of water body types is extremely rich, including large plateau lakes, inland rivers, glacial meltwaters, and mountain valleys. This high degree of diversity in space and water body types ensures that a model trained on this dataset can learn universal features of water bodies and possess good generalization potential.
The target domain (Dataset B) focuses on a highly challenging geographical unit—the Zhada Tulin (Zhada Earth Forest) in the Ali Prefecture of Tibet, China (Fig. 1). The imagery for this region is sourced from the Gaofen-2 (GF-2) satellite. Its resolution is consistent with that of GF-7, ensuring scale consistency for cross-dataset comparison. Located in the upper reaches of the Yarlung Zangbo River, the Zhada Tulin is a typical high-altitude arid region, and its hydrological environment differs significantly from the source domain: (1) Unique Morphology: The water systems flow through narrow gullies within the unique "earth forest" landscape, exhibiting irregular shapes and significant seasonal variations. (2) High Turbidity: The rivers carry large amounts of sediment, causing their spectral characteristics to differ vastly from the clear water bodies in the source domain. (3) Complex Background: The water bodies are highly coupled with the surrounding exposed sediments, both geomorphologically and spectrally, which can easily confuse segmentation algorithms. The unique geographical environment, hydrological features, and different sensor origins of the target domain collectively constitute a significant domain shift, making it an ideal testbed for validating the effectiveness of the proposed transfer learning strategy.
@@ 2.2 Two-Stage Transfer Learning Framework
    Stage 1: Source-Domain Pre-training. The SegFormer model is thoroughly trained using the source domain dataset (Dataset A). The objective of this stage is to enable the model to learn universal spatial and textural features of water bodies, resulting in a foundational model (Model A1) with strong generalization capabilities. To accelerate convergence, the encoder component is initialized with ImageNet pre-trained weights.
@@
To implement this framework, all original remote sensing images must undergo a standardized sample construction process (Fig. 3). This process involves two main steps: sliding window cropping and data augmentation, designed to convert the raw data into patches suitable for model training.
@@
Data Augmentation. To enhance the model’s generalization ability and effectively mitigate overfitting, a series of data augmentation techniques (Fig. 3c-f) are applied exclusively during the training phase. The employed strategies include:
@@
Through this sample construction pipeline, the final training and validation sets for the source and target domains were obtained. Specifically, 3,875 valid 512$\times$512 pixel patches were cropped from 207 original GF-7 images for the source domain (Dataset A), while 180 patches were obtained from 20 original GF-2 images for the target domain (Dataset B). Subsequently, to ensure an objective and consistent evaluation, both datasets were randomly split into training and validation sets at a 9:1 ratio. It is noteworthy that the training set for the target domain B contains only 162 samples (180 $\times$ 0.9), which further highlights the "small-sample" challenge in this specific region and underscores the necessity of the proposed transfer learning strategy.
@@ 2.3 Semantic Segmentation Model
This study selected SegFormer [38] as the core semantic segmentation network for the precise extraction of water bodies from remote sensing imagery. SegFormer is a highly efficient model based on a pure Transformer architecture, striking an excellent balance between accuracy and efficiency, making it particularly suitable for processing high-resolution remote sensing images. The model primarily consists of a hierarchical Transformer encoder and a lightweight all-MLP (Multilayer Perceptron) decoder head, as shown in Fig. 4.
Figure 4: Architecture of SegFormer, including a hierarchical Transformer encoder and an all-MLP decoder [38].
The encoder of SegFormer employs the Mix Transformer (MiT) [5] series as its backbone; this study specifically uses the MiT-B5 version. Unlike the traditional Vision Transformer (ViT) [6], MiT achieves efficient multi-scale feature extraction. It can effectively extract feature representations at different scales from the input image, which is crucial for identifying water bodies of various sizes and shapes. Through an improved self-attention mechanism and a position-encoding-free design, MiT significantly enhances computational efficiency and adaptability to different input sizes. This design paradigm, which has seen successful application in remote sensing image processing [36], is well-suited for handling the complex spatial relationships within remote sensing images.
@@ 2.4 Model Training and Implementation
The implementation and training of all models in this study were conducted based on the PyTorch [26] deep learning framework within a unified hardware environment to ensure the comparability of the results.
To effectively guide the training process of the SegFormer model and address issues such as class imbalance (e.g., significantly more background pixels than water pixels) and ambiguous object boundaries in the water body segmentation task, this study employs a compound loss function. This loss is a linear combination of a weighted Cross-Entropy Loss and a Dice Loss [23].
@@
The Cross-Entropy Loss [45], a standard pixel-level classification loss for image segmentation, aims to minimize the difference between the predicted pixel class probability distribution and the ground truth. To further mitigate the imbalance between foreground (water) and background pixels, different weights were assigned to each class. Specifically, the weight for the background class was 0.2289, and for the water class, it was 0.7711. By assigning a higher loss weight to the less frequent water class, the model is guided to pay more attention to its correct identification. For our binary semantic segmentation task, the formula is:
@@
    Optimizer: The AdamW optimizer [20] was used with an initial learning rate set to $6\times 10^{-6}$ and a weight decay of 0.01.
@@ 3.1 Source Domain Model Performance
To ensure the effectiveness of the subsequent transfer learning, the performance of the foundational model A1, trained on the source domain dataset A, was first evaluated. The segmentation performance of Model A1 on the validation set of the source domain is shown in Table 1.
Table 1: Performance of the source domain model A1 on its validation set.
Class IoU (%) F1-score (%) Precision (%) Recall (%)
Background 88.63 93.97 94.26 93.69
Water 68.80 81.52 80.77 82.28
@@
As can be seen from Table 1, the foundational model A1 achieved good segmentation results on its validation set. For the critical "Water" class, the Intersection over Union (IoU) reached 68.80%, and the F1-score reached 81.52% (note: in binary semantic segmentation, the F1-score is equivalent to the Dice coefficient). The precision and recall were 80.77% and 82.28%, respectively. These metrics indicate that Model A1 has developed strong generalization capabilities, laying a solid foundation for the subsequent knowledge transfer to the target domain.

### v2

@@ Abstract
The Tibetan Plateau, known as the ’Asian Water Tower’, faces significant water security challenges due to its high sensitivity to climate change. Advancing Earth observation for sustainable water monitoring is thus essential for building climate resilience in this region. This study proposes a two-stage transfer learning strategy using the SegFormer model to overcome domain shift and data scarcity—key barriers in developing robust AI for climate-sensitive applications. After pre-training on a diverse source domain, our model was fine-tuned for the arid Zhada Tulin area. Experimental results show a substantial performance boost: the Intersection over Union (IoU) for water body segmentation surged from 25.50% (direct transfer) to 64.84%. This AI-driven accuracy is crucial for disaster risk reduction, particularly in monitoring flash flood-prone systems. More importantly, the high-precision map reveals a highly concentrated spatial distribution of water, with over 80% of the water area confined to less than 20% of the river channel length. This quantitative finding provides crucial evidence for understanding hydrological processes and designing targeted water management and climate adaptation strategies. Our work thus demonstrates an effective technical solution for monitoring arid plateau regions and contributes to advancing AI-powered Earth observation for disaster preparedness in critical transboundary river headwaters.
@@
Funding
The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.
@@ 1 Introduction
Surface water bodies are crucial natural elements, and accurately mapping their spatial distribution is vital for sustainable water resource management, climate resilience building, and disaster risk assessment (McFeeters 1996). High-resolution remote sensing imagery has become a primary data source for this task.The integration of artificial intelligence with Earth observation systems represents a significant advancement in our capacity to support climate resilience and disaster risk reduction efforts. Traditional methods like the Normalized Difference Water Index (NDWI) (McFeeters 1996; Xu 2006) and threshold segmentation (Otsu et al. 1975) often struggle in complex environments where water’s spectral features are similar to the background, frequently requiring support from other indices like NDVI (Rouse Jr et al. 1973) or SAVI (Huete 1988).
The advent of deep learning has revolutionised remote sensing image analysis,offering new pathways for AI-driven Earth observation systems that support sustainability goals (Goodfellow, Bengio, and Courville 2016; Zhu et al. 2017; Ma et al. 2019; Zhang, Zhang, and Du 2016). Convolutional Neural Networks (CNNs) (LeCun et al. 2002), from early Fully Convolutional Networks (FCNs) (Long et al. 2015) to advanced architectures like U-Net (Ronneberger, Fischer, and Brox 2015), SegNet (Badrinarayanan, Kendall, and Cipolla 2017), and DeepLab (Chen et al. 2017, 2018), have significantly improved segmentation accuracy. More recently, Transformer-based models such as the Swin Transformer (Liu et al. 2021) and SegFormer (Xie et al. 2021), leveraging the global modelling power of the attention mechanism (Vaswani et al. 2017), have demonstrated outstanding performance. These methods, along with other innovations like multi-attention networks (Li et al. 2021), non-local networks (Wang et al. 2018), and hybrid Transformer-CNNs (Zhang et al. 2022), have greatly enhanced the automatic identification of ground features.
Despite these technological advancements, applying deep learning to real-world remote sensing is challenged by pressing geoscientific needs, particularly in the context of climate change. The Tibetan Plateau, for instance, is warming at approximately twice the global average rate (Q et al. 2011; Chen 2015), triggering significant hydrological changes like accelerated glacier melt and altered river runoff.These changes increase the frequency of extreme weather events, including flash floods and droughts, making accurate water body monitoring essential for disaster risk reduction and climate adaptation strategies. Traditional sparse monitoring stations are insufficient for capturing these dynamics. High-resolution remote sensing thus offers an indispensable tool for large-scale monitoring, which is crucial for understanding the plateau’s evolving water cycle and assessing its sustainability.
This study focuses on a particularly challenging yet critical region: the Zhada Tulin (Zhada Earth Forest) in the Ali Prefecture of Tibet. As a typical arid landscape in the northwestern Tibetan Plateau (Ding et al. 2006), its water systems are sensitive indicators of regional climate shifts. The area’s Xiangquan River is a vital headwater for a transboundary river system, directly impacting water security in downstream South Asia. Moreover, the unique ’earth forest’ geomorphology, a product of Quaternary tectonic uplift and climate change, makes it a natural laboratory for studying the plateau’s environmental evolution. Given the region’s vulnerability to climate-related hazards and its importance for regional water security, developing robust monitoring capabilities is crucial for both scientific understanding and disaster preparedness.
However, the practical application of deep learning models in such specific regions is severely hampered by two core issues. First, these models rely on large, high-quality annotated datasets, which are costly and labour-intensive to produce for remote sensing imagery. Second, a model trained in one geographical area often suffers a sharp performance decline when applied to another due to variations in imaging conditions and landscape features—a phenomenon known as ”domain shift” (Pan and Yang 2010; Yosinski et al. 2014).
To address these challenges, transfer learning provides an effective solution. While the ’pre-train and fine-tune’ paradigm using ImageNet (Krizhevsky, Sutskever, and Hinton 2017) is standard, its direct application to remote sensing yields limited results due to the unique characteristics of satellite imagery (Tuia, Persello, and Bruzzone 2016). Therefore, designing transfer learning strategies tailored for remote sensing applications remains a key research area.
@@
Against this background, this study proposes a two-stage transfer learning strategy based on the SegFormer model. We first train a robust base model on a diverse source domain and then fine-tune it using a small, targeted dataset from the Zhada Tulin area. This ’general-to-specific’ approach aims to achieve high-precision water body segmentation under small-sample and domain-shift conditions. This research not only provides a technical solution for water information extraction in the challenging Zhada region but also contributes to advancing Earth observation capabilities for climate resilience and sustainable water management in climate-sensitive environments.
@@ 2.1 Study Area and Datasets
The datasets for this study consist of a diverse source domain (Dataset A) designed for robust pre-training and a highly challenging target domain (Dataset B) for evaluating the model’s adaptability.
@@
Source Domain (Dataset A). The source domain was constructed to provide rich and generalizable learning samples. The imagery is primarily sourced from China’s Gaofen-7 (GF-7) satellite, whose sub-meter resolution (0.8-meter panchromatic and 3.2-meter multispectral) provides an excellent data foundation for fine-grained feature recognition. The spatial coverage is extensive, focusing on diverse plateau and mountainous environments, including arid lakes in Xinjiang, salt lakes in Qinghai, plateau lakes in Tibet, and mountain lakes and rivers in the Rocky Mountains of the United States. This high degree of diversity in geography and water body types ensures that a model trained on this dataset learns universal features of water bodies, possessing strong generalisation potential.
Target Domain (Dataset B): The Zhada Tulin The target domain focuses on a unique and challenging geographical unit—the Zhada Tulin (Zhada Earth Forest) in the Ali Prefecture of Tibet, China (Fig. 1). The imagery for this region is sourced from the Gaofen-2 (GF-2) satellite, with a resolution consistent with GF-7, ensuring scale consistency for cross-dataset comparison. The Zhada Tulin is an ideal testbed for our transfer learning strategy due to its distinct geoscientific significance and complex remote sensing characteristics.
  • Hydrological and Geological Features: Located north of the Bangong-Nujiang Suture Zone, the Zhada area features complex geological structures, primarily composed of Pliocene-Quaternary lacustrine and fluvial sediments (Zhu et al. 2007). Situated at an altitude of 3500-4200m, it experiences a high-altitude temperate arid climate with annual precipitation under 200mm and evaporation exceeding 2000mm, making water resources extremely scarce. The water systems mainly consist of the Xiangquan River, its tributaries, seasonal gully runoff, and a few spring outlets (Fang et al. 2003). Despite their small scale, these water bodies serve critical ecological functions: they create oasis effects that support unique plant communities (e.g., Tamarix, Salix) and sustain regional biodiversity (Q et al. 2023), while also playing a vital role in soil conservation against wind erosion.
  • Climatic Response and Research Value: Over the past three decades, the hydrology of the Zhada region has undergone significant changes in response to global warming, a trend consistent with broader observations across High Asia (W, van Beek L P H, and P 2010; F et al. 2014). These changes manifest as earlier snowmelt-driven runoff, an increased frequency of extreme events like flash floods, and the shrinkage of small lakes. Such dynamics make the Zhada Tulin an ideal case for studying the impact of climate change on arid plateau water cycles. Therefore, precise water body monitoring in this sensitive area is of great scientific value.
  • Challenges for Remote Sensing Segmentation: The unique environment of the Zhada Tulin presents significant challenges that constitute a domain shift from the source domain. These challenges include: (1) unique morphology, with water systems flowing through narrow, irregular gullies; (2) high turbidity, as rivers carry large amounts of sediment, altering their spectral characteristics; and (3) a complex background, where water bodies are spectrally and geomorphologically coupled with surrounding exposed sediments. These factors make the Zhada Tulin a perfect testbed for validating the effectiveness of the proposed transfer learning strategy.
@@ 2.2 Two-Stage Transfer Learning Framework
    Stage 1: Source-Domain Pre-training. The SegFormer model is thoroughly trained using the source domain dataset (Dataset A). The objective of this stage is to enable the model to learn universal spatial and textural features of water bodies, resulting in a foundational model (Model A1) with strong generalisation capabilities. To accelerate convergence, the encoder component is initialized with ImageNet pre-trained weights.
@@
To implement this framework, all original remote sensing images must undergo a standardised sample construction process (Fig. 3). This process involves two main steps: sliding window cropping and data augmentation, designed to convert the raw data into patches suitable for model training.
@@
Data Augmentation. To enhance the model’s generalisation ability and effectively mitigate overfitting, a series of data augmentation techniques (Fig. 3c-f) are applied exclusively during the training phase. The employed strategies include:
@@
Through this sample construction pipeline, the final training and validation sets for the source and target domains were obtained. Specifically, 3,875 valid 512$\times$512 pixel patches were cropped from 207 original GF-7 images for the source domain (Dataset A), while 180 patches were obtained from 20 original GF-2 images for the target domain (Dataset B). Subsequently, to ensure an objective and consistent evaluation, both datasets were randomly split into training and validation sets at a 9:1 ratio. It is noteworthy that the training set for the target domain B contains only 162 samples (180 $\times$ 0.9), which further highlights the ”small-sample” challenge in this specific region and underscores the necessity of the proposed transfer learning strategy.
@@ 2.3 Semantic Segmentation Model
This study selected SegFormer (Xie et al. 2021) as the core semantic segmentation network for the precise extraction of water bodies from remote sensing imagery. SegFormer is a highly efficient model based on a pure Transformer architecture, striking an excellent balance between accuracy and efficiency, making it particularly suitable for processing high-resolution remote sensing images. The model primarily consists of a hierarchical Transformer encoder and a lightweight all-MLP (Multilayer Perceptron) decoder head, as shown in Fig. 4.
Figure 4: Architecture of SegFormer, including a hierarchical Transformer encoder and an all-MLP decoder (Xie et al. 2021).
The encoder of SegFormer employs the Mix Transformer (MiT) (Cui et al. 2022) series as its backbone; this study specifically uses the MiT-B5 version. Unlike the traditional Vision Transformer (ViT) (Dosovitskiy et al. 2020), MiT achieves efficient multi-scale feature extraction. It can effectively extract feature representations at different scales from the input image, which is crucial for identifying water bodies of various sizes and shapes. Through an improved self-attention mechanism and a position-encoding-free design, MiT significantly enhances computational efficiency and adaptability to different input sizes. This design paradigm, which has seen successful application in remote sensing image processing (Wang et al. 2022), is well-suited for handling the complex spatial relationships within remote sensing images.
@@ 2.4 Model Training and Implementation
The implementation and training of all models in this study were conducted based on the PyTorch (Paszke 2019) deep learning framework within a unified hardware environment to ensure the comparability of the results.
To effectively guide the training process of the SegFormer model and address issues such as class imbalance (e.g., significantly more background pixels than water pixels) and ambiguous object boundaries in the water body segmentation task, this study employs a compound loss function. This loss is a linear combination of a weighted Cross-Entropy Loss and a Dice Loss (Milletari, Navab, and Ahmadi 2016).
@@
The Cross-Entropy Loss (Zhang and Sabuncu 2018), a standard pixel-level classification loss for image segmentation, aims to minimize the difference between the predicted pixel class probability distribution and the ground truth. To further mitigate the imbalance between foreground (water) and background pixels, different weights were assigned to each class. Specifically, the weight for the background class was 0.2289, and for the water class, it was 0.7711. By assigning a higher loss weight to the less frequent water class, the model is guided to pay more attention to its correct identification. For our binary semantic segmentation task, the formula is:
@@
    Optimizer: The AdamW optimizer (Loshchilov and Hutter 2017) was used with an initial learning rate set to $6\times 10^{-6}$ and a weight decay of 0.01.
@@ 3.1 Source Domain Model Performance
To ensure the effectiveness of the subsequent transfer learning, the performance of the foundational model A1, trained on the source domain dataset A, was first evaluated. The segmentation performance of Model A1 on the validation set of the source domain is shown in Table 3.1.
@@
Performance of the source domain model A1 on its validation set. Class IoU (%) F1-score (%) Precision (%) Recall (%) Background 88.63 93.97 94.26 93.69 Water 68.80 81.52 80.77 82.28
As can be seen from Table 3.1, the foundational model A1 achieved good segmentation results on its validation set. For the critical ”Water” class, the Intersection over Union (IoU) reached 68.80%, and the F1-score reached 81.52% (note: in binary semantic segmentation, the F1-score is equivalent to the Dice coefficient). The precision and recall were 80.77% and 82.28%, respectively. These metrics indicate that Model A1 has developed strong generalisation capabilities, laying a solid foundation for the subsequent knowledge transfer to the target domain.